Y= a X + b is the equation of a straight line. A student plot X vs Y instead of Y vs X, what will be the slope and the intercept on his graph in terms of a and b.
Answer #1

Y = aX + b

so if we solve for X

X = Y/a - b/a

so, on his graph y = (1/a) x +(-b/a)

slope m = 1/a

c = -b/a      
Answer #2

so if we solve for X

a X = Y - b

X = 1/a* Y - b/a

so slope is 1/a

and intercept is -b/a
Answer #3

Y = aX + b


so.. X = (1/a)Y + ( -b/a)


so.. slope = 1/a

intercept = -b/a
Answer #4

Y - b = aX

X = Y/a - b/a

so, on his graph y = (1/a) x +(-b/a)

slope m = 1/a

c = -b/a